d the peak inverse voltage on diode if the peak voltage for half wave rectifier circuit is 20V and diode cut in voltage is 0.3 will be
1
Expert's answer
2020-11-10T10:01:17-0500

PIV is the maximum reverse bias voltage that can appear on a diode in a given circuit if the rated power of the PIV is less than this diode breakdown value. For a rectifier, "PIV=V_m-V_d=20-0. 3=19.7" V.

Explanation taken from the site: https://www.sanfoundry.com/electronic-devices-circuits-questions-answers-half-wave-rectifier/ (Electronic Devices and Circuits Questions and Answers – Half-Wave Rectifier; task 7)